Ques1) Two square metal plates A and B  are of the same thickness and material. The side of B is twice of that A. These are connected in series. Let the resistances of A and B are denoted by RA and RB. Then show that RA / RB =1.

Grade:12

2 Answers

himanshu dev sharma
14 years ago
Ramesh V
14 years ago

R = pL/A

R1 = pL/L*t 

and R2 = p(2L)/2L*t  = pL/L*t 
which gives : R1 / R2 = 1

--

Please feel free to post as many doubts on our disucssion forum as you can. If you find any question difficult to understand - post it here and we will get you the answer and detailed solution very quickly. We are all IITians and here to help you in your IIT JEE preparation. All the best.

Regards,
Naga Ramesh
IIT Kgp - 2005 batch